The maximum wavelength that an electromagnetic wave can have and still eject electrons from a metal surface is 528 nm. What is the work function W0 of this metal? Express your answer in electron volts.
Answer #1

As work function,

W0 = hf0

W0 = hc/\lambda0

W0 = (6.64*10-34)(3*108)/(528*10-9)

W0 = 0.0377*10-17 J

W0 = 3.77*10-19 J

W0 = 2.356 eV